if the perimeter of a rectangle is p m and its area is a sq.m then what is the relation between them.
Answer by Alan3354(69443)

if the perimeter of a rectangle is p m and its area is a sq.m then what is the relation between them.
-------------
p = 2L + 2W
a = L*W --> L = W/a
---------
p = 2W/a + 2W
p = 2W(a+1)/a
------
Similar manipulations can be done, but none of the results are useful.
p can't be expressed in terms of "a" without including either L or W.